# Global Prediction of COVID-19 Variant Emergence Using Dynamics-Informed Graph Neural Networks

### v2

@@ Abstract
During the COVID-19 pandemic, a major driver of new surges has been the emergence of new variants. When a new variant emerges in one or more countries, other nations monitor its spread in preparation for its potential arrival. The impact of the variant and the timing of epidemic peaks in a country highly depend on when the variant arrives. The current methods for predicting the spread of new variants rely on statistical modeling, however, these methods work only when the new variant has already arrived in the region of interest and has a significant prevalence. The question arises: Can we predict when (and if) a variant that exists elsewhere will arrive in a given country and reach a certain prevalence? We propose a variant-dynamics-informed Graph Neural Network (GNN) approach. First, We derive the dynamics of variant prevalence across pairs of regions (countries) that applies to a large class of epidemic models. The dynamics suggest that ratios of variant proportions lead to simpler patterns. Therefore, we use ratios of variant proportions along with some parameters estimated from the dynamics as features in a GNN. We develop a benchmarking tool to evaluate variant emergence prediction over 87 countries and 36 variants. We leverage this tool to compare our GNN-based approach against our dynamics-only model and a number of machine learning models. Results show that the proposed dynamics-informed GNN method retrospectively outperforms all the baselines, including the currently pervasive framework of Physics-Informed Neural Networks (PINNs) that incorporates the dynamics in the loss function.
@@
Graph Neural Networks, COVID-19 Variants, Forecasting
@@ I Introduction
The COVID-19 pandemic presented an unprecedented global health crisis, severely affecting millions of people worldwide and demanding swift and effective responses from governments and healthcare providers [WHO]. As the pandemic progressed, multiple variants of COVID-19 emerged, each possessing unique genetic mutations that can significantly impact transmissibility, virulence, and even vaccine efficacy. During much of the COVID-19 epidemic, the surge in cases and severe outcomes (hospitalizations and deaths) have been driven by the emergence of new variants [Wiemken2023]. Consequently, monitoring the emergence and spread of these variants is crucial for devising appropriate public health measures and optimizing containment strategies [WHO2]. A critical factor driving the surge in a given region is the time of arrival of the new variant [Markov2023]. We can observe how a new variant, that has not appeared in region $A$ yet, spreads in region $B$. We can study the spread in the region $B$ to understand the properties of this new variant. If it is a highly transmissible or immune-evading variant [lambrou2022genomic], we can expect it to spread in the region $A$ eventually. However, when it happens in region $A$ precisely remains unknown. Therefore, making a good prediction of arrival time would lead to a more effective preparation and resource management.
@@
In this paper, we focus on the problem of predicting arrival of a new variant in a given region provided that it has appeared somewhere else. This can be posed as a set of two problems: (1) A new variant $X$ has arrived and is spreading in region $A_{1},A_{2},\dots$, but not yet in region $B$. Will it arrive in region $B$? (2) If yes, when will it arrive in region $B$? Since there may be some false positives and noise in the data, a variant may seem to have appeared but not spreading. Furthermore, a variant may start to spread, but quickly get dominated by another variant before it reaches a significant proportion of the circulating cases. This can be seen in Figure 1, which shows the proportions of a few different variants in the UK and Sweden. Therefore, we reformulate the two problems as the following:
@@
Figure 1: Plots showing the prevalence of a few different variants of COVID-19 over time in the UK and Sweden. The beginning of the spread of new variants can differ across countries by several weeks. This can be seen when examining the ”20I.Alpha.V1” and ”21J.Delta” variants in the provided plots, which show their appearance in the UK several weeks before reaching Sweden.
Problem 1 (Emergence Prediction).
Given the prevalence (proportion) of a variant in regions $A_{1},A_{2},\dots$, predict if the variant will ever reach a proportion $\theta$ (we say, dominance threshold) in region $B$.
Problem 2 (Delay Prediction).
@@
Given the prevalence (proportion) of a variant in regions $A_{1},A_{2},\dots$, predict when the variant will reach a proportion of $\theta$ in region $B$.
Here, the term “proportion” refers to the fraction of cases created by the variant under consideration of the total cases.
@@
It naturally follows from the above description that Problem 1 is a classification problem while Problem 2 is a regression problem. In practical scenarios, we envision addressing Problem 1 first, and only tackling Problem 2 if the outcome is positive. This is also important for the sake of evaluation as we can answer Problem 2 only if the variant actually arrives in the region of interest. Existing work has focused on the problem of predicting the prevalence of a variant when it has already arrived in the region of interest [sun2020forecasting, hu2020early, wu2020nowcasting]. Specifically, logistic regression turns out to provide a good estimate of the prevalence over time [lr01, lr02]. However, these techniques require the variant to have a non-zero proportion already in the region of interest. To the best of our knowledge, no prior work exists predicting the delay to reach a certain prevalence even when there is zero prevalence in the region. We show that Problem 1 (Emergence prediction) is a simpler problem (albeit, non-trivial) that can be solved by a decision tree. However, Problem 2 (delay detection) is more difficult to solve and requires more sophisticated techniques.
@@
When a new variant emerges with evolutionary favorable properties, the only way it can be transferred to a different region is through a host (human). This encourages the idea of using an underlying network of mobility to address the proposed problems. Since reaching a certain prevalence may depend on the other currently circulating variants, their dynamics (how fast one variant can spread over others) also play a role. Therefore, we propose a variant dynamics-informed Graph Neural Network (GNN) that utilizes a network of mobility and features inspired by variant dynamics to solve the proposed problems. This is different than typical Physics-Informed Neural Networks (PINNs) where the dynamics act as a regularization for the loss [raissi2019physics]. Here, we show that our approach of constructing appropriate features results in lower errors than incorporating the dynamics in the loss function.
@@
Our key contributions represent a novel effort in addressing the challenge of predicting the emergence of COVID-19 variants at a global level, and in doing so establishing a new benchmark for evaluating Emergence Detection and Delay Prediction on the CoVariants dataset [covariants]. More specifically, our contributions are as follows:
@@
    We derive the variant dynamics, under some assumptions, that leads to a model of the delay for a variant reaching a certain prevalence between a pair of countries.This model serves as a baseline in our exploration, providing valuable insights that can be leveraged by our GNN and other Machine Learning (ML) models.
@@
    We develop novel adaptations of GNNs which incorporate disease dynamics into the models. This innovative approach significantly improves the predictive performance, by accounting for the complex inter-dependencies between countries. Through some experiments, we demonstrate that our approach leads to superior results compared to several ML methods, including the currently pervasive framework of PINNs that include the dynamics in the loss function.
  3. 3.
    We make our evaluation pipeline publicly available  [^1] GitHub Link so that it can be used by any user-defined PyTorch model for evaluation, thus providing a new benchmark for the community.
@@ II-A Variants and Variant Dynamics
In the context of infectious diseases, the term ”variant” refers to a version of a virus with some changes in its genetic material, known as genome [CDC_class]. These changes happen through genetic mutations and can affect the virus’s characteristics, like how easily it spreads, how severe the illness it causes is, and whether it can evade the immune system or not. Variants can be categorized in multiple ways based on genetic differences, and the significance of these differences can vary. Some categorizations focus on a common ancestor, while others may specifically highlight mutations in key regions of the virus’s genome [CDC_class]. In our study, we rely on the categorizations provided by the CoVariants dataset, discussed further in Section IV-A.
(a) 21J.Delta vs 20I.Alpha.V1
(b) 21K.Omicron vs 21J.Delta
Figure 2: Example semi-log plots of variant proportions from data in the CoVariants dataset.
Furthermore, it’s worth noting that the proportions of different variants can be analyzed over time. Some studies have observed that these proportions follow a straight line when plotted on a semi-logarithmic scale. Therefore, they leverage this fact to understand the dynamics of variant emergence and prevalence in populations [Beesley2023]. We also observe this in our data as shown in Figure 2. However, to the best of our knowledge, existing studies do not derive any dynamics capable of predicting delays in the emergence of variants between multiple regions/countries.
@@ II-B Graph Neural Networks and Epidemics
In response to the COVID-19 pandemic, many efforts were made to develop forecasting models capable of predicting the progression and future trends of the COVID-19 pandemic [US, EU, DU]. However, these forecasting approaches have primarily focused on predicting overall case counts, hospitalizations, or mortality rates, rather than predicting delays in variant emergence. The ability to predict the timing of variant-specific outbreaks can significantly enhance public health preparedness and response strategies, allowing for timely interventions such as targeted vaccination campaigns and tailored public health measures [WHO2].
Many Deep Learning (DL) techniques have been explored for predicting the dynamics of infectious diseases. While many models use methods like GNNs [scarselli2008graph], and PINNs [pinnraissi2019physics] for forecasting, some approaches have also focused on genetic sequence modeling for variants [Nagpal2022, Basu2021].
E. Kharazmi et al. [pinnkharazmi2021identifiability], use PINNs in epidemiological models to better capture the dynamics through the underlying Ordinary Differential Equations (ODEs). These works are normally variations of the classical Susceptible-Infectious-Removed (SIR) model with additional compartments [pinnkiselev2023delay]. Extending this, A. Rodríguez et al. [pinnrodriguez2023einns] propose a framework for forecasting, Epidemiologically-informed Neural Networks (EINNs) that integrates epidemiological knowledge embedded in ODE models into neural networks. The core idea is utilizing the learned representations from a PINN using transfer learning along with a neural network that ingests multi-modal data. Other extensions include Epi-DNNs from X. Ning et al. [pinnning2023epi], in which the neural network expresses the unknown parameters in the compartmental model, and the ODEs are solved using the Runge-Kutta method.
Since infectious diseases inherently present across multiple regions over a period of time, network models, like GNNs help in capturing the low-level graph effects. Works [GTNN, STAN, transfergraph, Sung2022, pinnganesan2021spatio], including Gao et al. [STAN] rely on GNNs to predict the total number of cases for a fixed period in the future using a spatio-temporal attention network. These models typically map locations (e.g., a county or a state) to nodes on a graph and construct the edges based on geographical proximity and demographic similarity between locations. As an extension, [pinnseo2019physics, pinnseo2019differentiable] combines spatio-temporal GNNs with PINNs. [chopra2022differentiable, chopra2021deepabm] use agent-based models on top of these to further capture low-level interactions between populations in different compartments.
While some works do focus on using the spatio-temporal networks with location aware features [pinndeng2019graph], including mobility data [pinnkapoor2020examining], to the best of our knowledge, no studies have attempted to predict the emergence of a variant in one country based on its occurrence in another (Problems 1 & 2), through variant informed dynamics.
In summary, instead of predicting the spread of the virus (reported cases, deaths, and hospitalizations), as others have done, we aim to predict if and when a new variant will emerge in a region given that it has emerged elsewhere. This involves deriving the dynamics of variant spread and incorporating these insights into a GNN. Notably, our strategy revolves around crafting relevant features rather than altering the loss function, providing a distinctive and effective method for predicting variant emergence.
@@ III-A Variant Dynamics
First, we develop an understanding of the dynamics that play a role in the spread of a new variant and how we can utilize them to make informed predictions. These dynamics can be epidemiological, spatial, or temporal. We start by extracting each variant’s global growth rates (see Section III-B) from the data. This value provides an indication of the dominance of the variants over other competing ones. Next, we build a model based on the dynamics of infectious diseases and the derived growth rates. This model is used as a baseline for delay prediction. Finally, we consider different approaches in which we can incorporate some disease dynamics using GNNs that are chosen for their ability to intuitively model spatial dynamics, with the goal of enhancing the prediction accuracy of our model. Throughout the text, we use the terms ”prevalence” and ”prevalence ratio” as defined below:
$\text{Prevalence}(p_{i})=\frac{\text{\# of COVID cases of a given variant }i}{\text{Total \# of all cases}}$ (1)
Prevalence $\text{Prevalence Ratio}(r_{ij})=\frac{p_{i}}{p_{j}}$ (2)
where $p_{i}$ is the prevalence of the variant $i$, and $p_{j}$ is the prevalence of the most prevalent variant, $j$, that is not $i$. For a variant to be considered dominant, it must reach an $r_{ij}$ value of at least $\frac{1}{3}$. We define this value as the dominance threshold, $\theta$.
@@ III-B Variant Growth Rates
We first derive the dynamics of two competing variants in one region. While logistic regression has been widely used to fit these dynamics, for completeness, we present a derivation. We start by using a model of infectious diseases that generalizes a large class of models [SIkJalpha]:
@@
The above further assumes that the mean serial intervals of the two variants are close, i.e., $b_{i}=b_{j}=b$. Note that we assumed full cross-immunity which is true for a number of variants before the Omicron sub-variants [Chen2023]. If there is only a partial cross immunity, $\bar{S}_{i}(t)\neq\bar{S}_{j}(t)$. However, in a short interval, the new immunity (reduction is susceptibility) created by the variants is small. As a result, it can be shown that $\bar{S}_{i}(t)\approx\beta_{ij}^{\prime}\bar{S}_{j}(t)$ for some constant $\beta_{ij}^{\prime}$ which can be absorbed in the parameter $\beta_{ij}$ above. Solving the above recurrence relation, we get:
@@
The parameter $S_{ij}$ can be referred to as the relative growth rate of variant $i$ over variant $j$ [SIkJalpha]. Now, consider a scenario in two regions $A$ and $B$$A$ has the variant $i$ emerging over the previously dominant variant $j$; but variant $j$ never reached region $B$ where variant $k$ is dominant. This makes it difficult to assess the potential impact of variant $i$ on region $B$ when we only know $S_{ij}$. To deal with this challenge, we note that the parameter $S{ij}$ should ideally be independent of the region (this may not hold due to the simplifying assumptions above, so we actually have a region-specific relative growth rate $S^{(p)}_{ij}$ for region $p$). Therefore, we can attempt to find the growth-rate of any variant $i$ with a fixed variant 00, specifically, the original COVID-19 variant. For ease of notation, we represent the growth advantage of variant $i$ with respect variant 00 as $S_{i}$, with $S_{0}=0$, and refer to these values as global growth rates. Ideally, $S_{ij}=S_{i}-S_{j}$, irrespective of the region where the $S_{ij}$ is estimated. To estimate these the global growth rates for all the variants that exist by the time $t$ over all regions $p$, we can set up the following system of linear equations:
@@
where $S^{(p)}_{ij}$ is the growth advantage of variant $i$ over variant $j$ in region $p$ and $\epsilon_{ijp}$ are error terms. The solution is given by minimizing the sum of squares of $\epsilon_{ijp}$. We, then, define an objective function and solve for the $S_{i}$ value as
@@
These growth rates are used in our proposed methods for both Emergence Prediction and Delay Prediction. We note in our Benchmarking study, Section IV-E, that existing simple ML models with our constructed features produce very high accuracy for the former problem. Therefore, we focus on our approaches for the latter problem next.

### v3

@@ Abstract
During the COVID-19 pandemic, a major driver of new surges has been the emergence of new variants. When a new variant emerges in one or more countries, other nations monitor its spread in preparation for its potential arrival. The impact of the new variant and the timings of epidemic peaks in a country highly depend on when the variant arrives. The current methods for predicting the spread of new variants rely on statistical modeling, however, these methods work only when the new variant has already arrived in the region of interest and has a significant prevalence. Can we predict when a variant existing elsewhere will arrive in a given region? To address this question, we propose a variant-dynamics-informed Graph Neural Network (GNN) approach. First, we derive the dynamics of variant prevalence across pairs of regions (countries) that apply to a large class of epidemic models. The dynamics motivate the introduction of certain features in the GNN. We demonstrate that our proposed dynamics-informed GNN outperforms all the baselines, including the currently pervasive framework of Physics-Informed Neural Networks (PINNs). To advance research in this area, we introduce a benchmarking tool to assess a user-defined model’s prediction performance across 87 countries and 36 variants.
@@
Epidemiology, Graph Neural Networks, Public Health Planning, Forecasting
@@ I Introduction
The COVID-19 pandemic presented an unprecedented global health crisis, severely affecting millions of people worldwide and demanding swift and effective responses from governments and healthcare providers [1]. As the pandemic progressed, multiple variants of COVID-19 emerged, each possessing genetic mutations that can significantly impact transmissibility, virulence, and even vaccine efficacy. During much of the COVID-19 epidemic, the surge in cases and severe outcomes (hospitalizations and deaths) have been driven by the emergence of new variants [2]. Consequently, monitoring the emergence and spread of these variants is crucial for devising appropriate public health measures and optimizing containment strategies [3]. A critical factor driving the surge in a given region is the arrival time of the new variant [4]. We can observe how a new variant, that has not appeared in region $A$ yet, spreads in region $B$. We can study the spread in the region $B$ to understand the properties of this new variant. If it is a highly transmissible or immune-evading variant [5], we can expect it to spread in the region $A$ eventually. However, precisely when it would happen in region $A$ remains unknown. Making a good prediction of arrival time would lead to more effective preparation and resource management.
We focus on the problem of predicting the arrival of a new variant in a given region provided that it has appeared somewhere else. Due to testing delays and the fact that not all cases are genomically analyzed, it is difficult to pin down when a variant arrives. Therefore, we consider measuring the delay to reach a certain proportion of prevalence.
@@
Figure 1 shows the proportions of different variants over time in the United Kingdom and Sweden. We can see that there is often a clear delay (see 20.Alpha.V1 and 21.J.Delta) in arrival of the variants in Sweden. Furthermore, a variant may start to spread, but quickly get dominated by another variant before it reaches a significant proportion of the circulating cases (e.g., 21.A.Delta in Figure 1). Therefore, we reformulate the problem as the following:
@@
Figure 1: Plots showing the prevalence of a few different variants of COVID-19 over time in the UK and Sweden. The beginning of the spread of new variants can differ across countries by several weeks. This can be seen when examining “20I.Alpha.V1” and “21J.Delta” variants in the provided plots, showing their appearance in the UK several weeks before reaching Sweden.
Problem 1 (Emergence Delay Prediction).
@@
Given the prevalence (proportion) of a variant in regions $A_{1},A_{2},\dots$, predict when the variant will reach a proportion of $\theta$ in region $B$, provided that this variant has not yet reached region $B$.
Here, the term “proportion” refers to the fraction of new cases created by the variant under consideration out of total new cases. Note that this is a regression problem – we seek to predict a quantity, which is the delay between the current date and the date of reaching the desired threshold. Further, we wish to perform this prediction for a given region before any sample of the variant of interest is observed in that region.
@@
Prior works have focused on the problem of predicting the prevalence of a variant when it has already arrived in the region of interest [6, 7]. Specifically, logistic regression turns out to provide a good estimate of the prevalence over time [8, 9]. However, these techniques require the variant to have a non-zero proportion already in the region of interest. To the best of our knowledge, no prior work exists predicting when the delay reaches a certain prevalence even when there is zero prevalence in the region.
@@
When a new variant emerges with evolutionarily favorable properties, it can only be transferred to a different region through a host (in the case of COVID-19 – a human). This encourages the idea of using an underlying network of mobility to address the proposed problem. Since reaching a certain prevalence may depend on other currently circulating variants, their dynamics (how fast one variant can spread over others) also play a role. Therefore, we propose a variant dynamics-informed Graph Neural Network (GNN) that utilizes a network of mobility and features inspired by variant dynamics to solve the proposed problem. This is different than typical Physics-Informed Neural Networks (PINNs) where dynamics act as a regularization for the loss [10]. Here, we show that our approach of constructing appropriate features results in lower errors compared to the PINNs approach.
@@
Our key contributions represent a novel effort in addressing the challenge of predicting the emergence of COVID-19 variants at a global level, and in doing so establishing a new benchmark for evaluating Delay Prediction on the “CoVariants” dataset [11]. We expect that this benchmark will help build the capacity to predict arrival times of the future variants of COVID-19 and other outbreaks. More specifically, our contributions are as follows:
@@
    We develop novel adaptations of GNNs that account for complex inter-dependencies between countries using GNNs while incorporating variant delay dynamics at the node level. To the best of our knowledge, we are the first to derive these delay dynamics for variants. Experiments demonstrate that our approach leads to superior results compared to several Machine Learning (ML) methods, including the currently pervasive framework of PINNs that integrate dynamics in the loss function.
@@
    We make our evaluation pipeline publicly available so that it can be used to evaluate any user-defined PyTorch model [^1] https://github.com/itssmutnuri/gnnvariants. This will advance public health research and enable public health experts and policymakers to leverage this benchmark pipeline for improvement of their models and consequently, for enhancing global health outcomes.
@@ II-A Variants and Variant Dynamics
In the context of infectious diseases, the term “variant” refers to a version of a virus with some changes in its genetic material, known as genome [12]. These changes happen through genetic mutations and can affect the virus’s characteristics, like how easily it spreads, how severe the illness it causes is, and whether it can evade the immune system or not. Variants can be categorized in multiple ways based on genetic differences, and the significance of these differences can vary. Some categorizations focus on a common ancestor, while others may specifically highlight mutations in key regions of the virus’s genome [12]. In our study, we rely on the categorizations provided by the CoVariants dataset [11].
Figure 2: Sample semi-log plots of variant proportions from data in the CoVariants dataset.
Furthermore, the proportions of different variants can be analyzed over time. Some studies have observed that these proportions follow a straight line when plotted on a semi-logarithmic scale. They leverage this fact to understand the dynamics of variant prevalence in populations [13]. We also observe this in our data as shown in Figure 2. However, to the best of our knowledge, existing studies do not derive any dynamics capable of predicting delays in the emergence of variants between multiple regions/countries.
@@ II-B Graph Neural Networks and Epidemics
Many efforts were made to develop forecasting models capable of predicting the progression and future trends of the COVID-19 pandemic [14, 15, 16]. However, these forecasting approaches have primarily focused on predicting the spread of the virus rather than predicting delays in variant emergence. Predicting the timing of variant-specific outbreaks can significantly enhance public health preparedness and response strategies, allowing for timely interventions such as targeted vaccination campaigns and tailored public health measures [3].
Many Deep Learning (DL) techniques have been explored for this purpose, including GNNs which help capture the graph effects of the spread of infectious diseases over time across regions [17, 18, 19, 20, 21]. Other methods like PINNs  [22, 23, 24], agent-based models [25, 26], and a combination of spatio-temporal networks with location-aware features [27], including mobility data [28] have also been used.
Instead of only predicting virus spread, as others have done, we aim to predict when a new variant will emerge in a region given that it has emerged elsewhere. This involves deriving dynamics of variant spread and integrating these insights into a GNN. Notably, our strategy revolves around crafting relevant features rather than altering the loss function, providing a distinctive and effective approach.
@@ III-A Variant Dynamics
First, we develop an understanding of the dynamics that play a role in the spread of a new variant and how we can utilize them to make informed predictions. We start by extracting each variant’s global growth rate from observed data. This value provides an indication of the dominance of the variants over other competing ones. Subsequently, we construct linear models based on the dynamics of infectious diseases and the derived growth rates to predict delays between regions, establishing this Dynamics-based Model as our baseline for Delay Prediction. Finally, we consider different approaches in which we can incorporate some disease dynamics using GNNs that are chosen for their ability to intuitively model spatial dynamics, with the goal of enhancing the prediction accuracy of our model.
Throughout the text, we use the terms “prevalence” and “prevalence ratio”, where the “prevalence”, $p$, is the proportion of cases of a given variant, and the “prevalence ratio” is defined as $r_{ij}=\frac{p_{i}}{p_{j}}$, where $p_{i}$ is the prevalence of the variant $i$, and $p_{j}$ is the prevalence of the most prevalent variant, $j$, that is not $i$. In our experiments, a variant is considered dominant if it reaches an $r_{ij}$ value of at least $\frac{1}{3}$. We define this value as the dominance threshold, $\theta$.
@@ III-B Variant Growth Rates
We first derive the dynamics of two competing variants in one region. While logistic regression has been widely used to fit these dynamics, for completeness, we present a derivation. We start by using a model of infectious diseases that generalizes a large class of models [29]:
@@
The above further assumes that the mean serial intervals of the two variants are close, i.e., $b_{i}=b_{j}=b$. Note that we assumed full cross-immunity which is true for a number of variants before the Omicron sub-variants [30]. If there is only a partial cross immunity, $\bar{S}_{i}(t)\neq\bar{S}_{j}(t)$. However, in a short interval, the new immunity (reduction is susceptibility) created by the variants is small. As a result, it can be shown that $\bar{S}_{i}(t)\approx\beta_{ij}^{\prime}\bar{S}_{j}(t)$ for some constant $\beta_{ij}^{\prime}$ which can be absorbed in the parameter $\beta_{ij}$ above. Solving the above recurrence relation, we get:
@@
Here, constant $b$ is the mean serial interval – the average time between two successive infections, and $\beta_{ij}$ is the ratio of transmission rates between the two competing variants. Parameter $S_{ij}$ can be referred to as the relative growth rate of variant $i$ over variant $j$ [29].
Now, consider a scenario in two regions $A$ and $B$$A$ has the variant $i$ emerging over the previously dominant variant $j$; but variant $j$ never reached region $B$ where variant $k$ is dominant. This makes it difficult to assess the potential impact of variant $i$ on region $B$ when we only know $S_{ij}$. To deal with this challenge, we note that the parameter $S{ij}$ should ideally be independent of the region (this may not hold due to the simplifying assumptions above, so we actually have a region-specific relative growth rate $S^{(p)}_{ij}$ for region $p$). Therefore, we can attempt to find the growth-rate of any variant $i$ with a fixed variant 00, specifically, the original COVID-19 variant. For ease of notation, we represent the growth advantage of variant $i$ with respect variant 00 as $S_{i}$, with $S_{0}=0$, and refer to these values as global growth rates. Ideally, $S_{ij}=S_{i}-S_{j}$, irrespective of the region where the $S_{ij}$ is estimated.
To estimate these the global growth rates for all the variants that exist by the time $t$ over all regions $p$, we can set up the following system of linear equations:
@@
where $S^{(p)}_{ij}$ is the growth advantage of variant $i$ over variant $j$ in region $p$ and $\epsilon_{ijp}$ are error terms. The solution is given by minimizing the sum of squares of $\epsilon_{ijp}$.
We, then, define an objective function and solve for the $S_{i}$ value as
@@
These growth rates are used in our proposed methods for Delay Prediction.